OpenStudy (anonymous):

Generate the first 5 terms of this sequence: f(1) = 1 and f(2) = 2, f(n) = f(n - 1) + f(n - 2), for n > 2

OpenStudy (anonymous):

@johnweldon1993 can you help /.\

OpenStudy (johnweldon1993):

of course /.\ lol

OpenStudy (anonymous):

Yays <3

OpenStudy (johnweldon1993):

f(1) = 1 f(2) = 2 f(n) = f(n-1) + f(n - 2) so for f(3) we have f(3) = f(3-1) + f(3 - 2) f(3) = f(2) + f(1) f(3) = 2 + 1 f(3) = 3 f(4) = f(4-1) + f(4 - 2) f(4) = f(3) + f(2) f(4) = 3 + 2 f(4) = 5 f(5) = f(5-1) + f(5 - 2) f(5) = f(4) + f(3) f(5) = 5 + 3 f(5) = 8 :)

OpenStudy (jdoe0001):

\(\large \begin{array}{ccllll} \textit{term number}&\textit{term value}\\ \hline\\ {\color{red}{ 1}}&1\\ {\color{red}{ 2}}&2\\ {\color{red}{ 3}}&f({\color{red}{ 3}}) = f({\color{red}{ 3}} - 1) + f({\color{red}{ 3}} - 2)\\ {\color{red}{ 4}}&f({\color{red}{ 4}}) = f({\color{red}{ 4}} - 1) + f({\color{red}{ 4}} - 2)\\ {\color{red}{ 5}}&f({\color{red}{ 5}}) = f({\color{red}{ 5}} - 1) + f({\color{red}{ 5}} - 2)\\ \end{array}\)

OpenStudy (anonymous):

1, 2, 3, 4, 5 1, 2, 3, 5, 7 1, 2, 3, 5, 8 1, 3, 5, 7, 9

OpenStudy (anonymous):

these are my choices

OpenStudy (johnweldon1993):

Mmhmm f(1) = 1 f(2) = 2 f(3) = 3 f(4) = 5 f(5) = 8

OpenStudy (anonymous):

thanksss darling ^.^